Which of these physical quantities has base units kgm2s-2?


A. Pressure B. Energy C. Power D. Momentum




Intensity is defined as power per unit area, What is the units of intensity?


1
Expert's answer
2021-08-19T14:19:37-0400

Pressure(P)="\frac{F}{A}"

"F=[M^1L^1T^{-2}]\\A=[M^0L^2T^0]"

"P=\frac{[M^1L^1T^{-2}]}{[M^0L^2T^0]}=[M^1L^{-1}T^{-2}]"

Pressure="N \/m^2"

Energy (E)=(w)="F\times dr"


"W=[M^1L^1T^{-2}][L^1]=[M^1L^2T^{-2}]"

Unit ="Kgm^2sec^{-2}"

Power="\frac{E}{t}=\frac{[ML^2T^{-2}]}{[T]}=[ML^2T^{-3}]"

Unit=J/sec

Momentum (p) =mv

"P=[M][LT^{-1}]=[MLT^{-1}]"

Unit =kgm/sec

Option (b) is correct option

Intensity(I)="\frac{P}{A}"

Power=J/sec

Area=m2

"I=\frac{J/sec}{m^2}=\frac{J}{m^2\times sec}"